# DyBBT: Dynamic Balance via Bandit inspired Targeting for Dialog Policy with Cognitive Dual-Systems

### v1

@@ 4.1 Experimental Setup
Datasets. We conduct experiments on two of the most prominent TODS benchmark datasets which are also used in baselines. The Microsoft Dialog Challenge platform Li et al. (2018); Zhao et al. (2024); Niu et al. (2024) for single domain, while the MultiWOZ2.1 dataset Budzianowski et al. (2018) for multi domains. Statistics in Appendix B.1.
@@
Evaluation Metrics. For single-domain tasks: success rate, average turns, and reward (following EIERL Zhao et al. (2025): +2t+2t for success, t-t for failure, 1-1 for every turn). For multi domain: Inform, Success, Book rates, and Avg. Turns (formulas in Appendix B.3).
Implementation Details. Following EIERL for fair comparison, dialogs are capped at 30 (single domain) and 40 (multi domain) turns. Training runs for 500 epochs (single) and 10K epochs (multi). DyBBT uses the same Qwen3 (0.6B–8B) for both S1 and S2. Full details in Appendix B.5.
4.2 Main Results
Table 1: Evaluation results for all agents across the three single domain datasets are provided, with the highest value in each metric column highlighted in bold. Epochs (50, 250, 500) represent early, mid, and post convergence training stages. Baselines sourced from Zhao et al. (2025).
[…]
4.2.1 Performance on Single Domain Tasks
The evaluation results on single domain dialog tasks are presented in Table 1. DyBBT demonstrates strong performance across all three domains. The results reveal that DyBBT’s cognitive enables more efficient policy learning: by dynamically allocating computational resources based on real-time cognitive signals, DyBBT achieves higher task success with significantly fewer dialog turns compared to methods relying on static exploration heuristics, population level evolution or GPT-4 based policy. This efficiency gain is particularly pronounced in complex domains like Taxi, where slot dependencies create challenging exploration landscapes that DyBBT navigates more effectively through its principled switching mechanism.
4.2.2 Performance on Multi Domain Task
Results on the challenging MultiWOZ dataset are provided in Table 4 (Appendix D.1). While EIERL’s success rate drops significantly in this complex multi domain setting, highlighting the scalability limits of its population based approach, DyBBT maintains strong performance. DyBBT-8B performs slightly better than AutoTOD/ProTOD (AutoTOD, ProTOD), and using GPT-4 as S2 yields SOTA results, showing that DyBBT matches strong LLM baselines while being more efficient. This is enabled by the structured cognitive state and dual-system design, which provide a domain agnostic inductive bias without requiring task specific tuning.
4.2.3 Training Efficiency and Convergence
Figure 3 illustrates the learning curves of DyBBT compared to baselines. DyBBT converges faster and achieves higher asymptotic performance across all domains, outperforming EIERL significantly even at epoch 50. This accelerated learning stems from the meta-controller’s active guidance of exploration from the outset, which systematically targets under explored or uncertain regions in $\mathcal{C}$ rather than relying on random exploration or high-variance evolutionary mechanisms.
Furthermore, DyBBT exhibits consistent scaling with model size, for instance, success rates improve from 80.34% to 89.52% in the single domain Movie task and from 78.2% to 84.1% in the multi domain setting when scaling from 0.6B to 8B parameters. This trend indicates that the dual-system architecture effectively harnesses the increased representational capacity of larger backbone models. When coupled with the meta-controller’s efficient resource allocation with Qwen3’s native switching mechanism in balancing performance and computational cost (Appendix D.6). DyBBT underscores its practical viability for real-world deployment.
@@ 4.2.3 Training Efficiency and Convergence
4.2.4 Summary of Strengths
The main results demonstrate that DyBBT achieves state-of-the-art performance through: Dynamic Exploration-Exploitation Balance: The meta-controller’s bandit inspired switching rule allows DyBBT to dynamically allocate expensive S2 reasoning only when necessary, leading to highly efficient exploration. Scalability with Model Size: DyBBT benefits predictably from larger backbone models, making it well suited for future advancements in LLM capabilities. Strong Generalization: Consistent performance across both single and multi domain tasks shows that the cognitive state representation captures universal dialog dynamics. Computational Practicality: Unlike population based methods (EIERL) or full GPT-4.0 approaches, DyBBT maintains moderate computational overhead during both training and inference.
@@ 4.3 Ablation Experiment
Ablation settings and results (Table 5) are detailed in Appendix D.2, revealing that: Meta-Controller is crucial. Removing it causes the most severe performance degradation, confirming its essential role in dynamically orchestrating the exploration-exploitation trade-off. Both conditions are necessary but asymmetric: Removing the confidence condition (CC) causes a more substantial performance drop than removing the exploration condition (EC), validating our hybrid design. This indicates that mitigating S1’s overconfidence is slightly more critical than targeted exploration for robust performance. Cognitive State design is vital. Replacing it with the raw belief state causes catastrophic performance collapse, confirming the necessity of our low dimensional, interpretable representation. While the learned alternative performs reasonably well, it still underperforms our hand-designed features, justifying our cognitively inspired approach. All state dimensions contribute meaningfully. Removing any single dimension causes noticeable performance degradation, with dialog progress ($d_{t}$) being the most impactful individual component, followed by user uncertainty ($u_{t}$) and slot dependency ($\rho_{t}$). Knowledge Distillation enables continuous improvement. Disabling it reduces final performance, confirming its role in facilitating long term efficiency gains through systematic learning from S2’s demonstrations.
4.4 Human Evaluation
@@
Automated metrics cannot fully capture the quality of decision making, such as action appropriateness or the timing of reasoning. We thus conducted a human evaluation (details in Appendix C) focusing on the meta-controller’s switching decisions. Ten NLP researchers evaluated 200 dialog states from MultiWOZ, comparing DyBBT against random switching and S1-only baselines. Annotators assessed action appropriateness (5 point Likert scale) and whether invoking S2 was justified (binary judgment). Results show that DyBBT’s actions are rated significantly more appropriate than both baselines. Moreover, its decisions to invoke S2 align substantially better with human judgment than random switching, providing qualitative evidence that our meta-controller effectively identifies when deliberation is warranted a key affordance often missed by heuristic approaches.
DyBBT validates a core tenet of ecological psychology in artificial dialog: that intelligent exploration emerges not from global heuristics or raw compute, but from perceiving and responding to local affordances. Our framework successfully translates this principle into a scalable, high performing policy proving that cognitively grounded design is key to efficient, robust dialog systems.
@@ 5 Analysis
Our experimental results demonstrate that DyBBT achieves state-of-the-art performance across multiple benchmarks. In this section, we analyze the underlying mechanisms that enable DyBBT’s effectiveness, providing insights into why and how our framework works.
5.1 The Emergent Structure of Cognitive State Space
@@
The cognitive state space $\mathcal{C}$ serves as the foundational bridge that enables the transfer of bandit exploration principles to the complex dialog POMDP. To empirically validate its utility, we analyze the visitation frequency of different regions within the discretized $\mathcal{C}$ over training (Fig. 5; detailed computation in Appendix B.5.3). The heatmap reveals a highly structured, non-uniform occupancy pattern, directly validating our core hypothesis. The meta-controller’s exploration is not random but strategically focused: in the early dialog phase ($d_{t}\in[0.0,0.2]$), it broadly explores across user uncertainty ($u_{t}$) for information gathering. In the mid-phase ($d_{t}\in[0.4,0.6]$), visitation concentrates in regions of medium-to-high $u_{t}$, targeting ambiguity resolution. In the late phase ($d_{t}>0.8$), activity focuses on states with low $u_{t}$, exploiting known information for task completion.
This phase dependent targeting demonstrates that $\mathcal{C}$ successfully captures the dialog’s dynamic “affordances”. The meta-controller learns to allocate its exploration budget to the most relevant regions of $\mathcal{C}$ for the current dialog stage, enabling highly efficient and context aware exploration. The effectiveness of $\mathcal{C}$ stems from its ability to distill the high dimensional belief state into a low dimensional, actionable representation, making principled exploration computationally feasible.
5.2 Adaptive Balancing Through Dual Triggers
The meta-controller’s hybrid triggering mechanism provides a robust solution to the exploration-exploitation dilemma by responding to different types of uncertainty:
Epistemic vs. Aleatoric Uncertainty Distinction: Two trigger conditions address fundamentally different types of uncertainty. The exploration condition ($n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}$) targets epistemic uncertainty, lack of knowledge about the environment that can be reduced through exploration. The confidence condition ($p_{t}^{S1}<\kappa$) addresses aleatoric uncertainty, inherent stochasticity or model limitations irreducible via exploration alone. Complementary Trigger Patterns: Analyzing 10,000 dialog turns reveals complementary triggering patterns (Fig. 5 in Appendix D.3). The exploration condition dominates in early training phases and for novel state regions, enabling systematic coverage of the state space. The confidence condition acts as a consistent safety net throughout training, preventing overreliance on a potentially flawed System 1. This complementary design ensures robustness across diverse dialog scenarios. Progressive Adaptation: The triggering rate evolves naturally with training progress. Initially, frequent System 2 invocations offer guided exploration and high quality demos. As training progresses and System 1 improves through distillation, the meta-controller automatically reduces System 2 usage, transitioning from guided exploration to autonomous operation. This adaptive balancing is key to DyBBT’s computational efficiency and crucially, it is the core manifestation of DyBBT’s ability to perceive and respond to the dynamic “affordances" of the dialog environment, ensuring the right cognitive system is invoked at the right time.
5.3 Knowledge Distillation as Implicit Policy Improvement
The knowledge distillation process creates a virtuous cycle that enables continuous policy improvement without additional environment interactions. The effectiveness of distillation is evidenced by the monotonic improvement of System 1 and corresponding reduction in System 2 invocation rate (Fig. 7 in Appendix D.3), demonstrating successful knowledge transfer.
@@ 5.4 Theoretical Intuitions and Empirical Alignment
Our theoretical analysis, though based on simplifying assumptions, is pragmatically validated by empirical results: Sublinear Regret as Validation of Core Assumptions. The empirical cumulative regret (Fig. 7) exhibits $\sqrt{T}$-like growth. This sublinear trend is not merely observational; it provides indirect empirical support for our key theoretical assumptions: The Lipschitz continuity of the reward in $\mathcal{C}$ (Assumption 1), and the approximate structure of MDP over $\mathcal{C}$ (Assumption 2). The alignment between theory and experiment suggests $\mathcal{C}$ effectively captures the latent structure enabling efficient exploration. Low Dimensional $\mathcal{C}$ Enables Practical Implementation. The consistent high performance of DyBBT using only a three dimensional cognitive state demonstrates that the essential features governing exploration (dialog progress, user uncertainty, slot dependency) can be distilled into a compact representation. This dimensionality reduction is theoretically motivated as the regret bound’s $\sqrt{\dim(\mathcal{C})}$ dependence (Appendix A.2.2).
5.5 Failure Mode Analysis and Limitations
Despite its strong performance, DyBBT exhibits three key failure modes that constrain its robustness. It is over reliant on cognitive state fidelity: the handcrafted $\mathbf{c}_{t}$ can misrepresent complex dialog dynamics such as abruptting intent shifts, leading the meta-controller to misjudge when to invoke System 2, either under exploring or wasting computation. It depends on high quality System 2 demonstrations: errors in System 2’s reasoning or selfevaluation, even if confident, can be distilled into System 1, corrupting its policy in a subtle cascading failure. It is sensitive to discretization: the heuristic 5 bins quantization of $\mathcal{C}$ can mask critical state variations, causing under exploration by treating strategically distinct states as identical. These limitations reveal a tension between DyBBT’s elegant, theory driven design and the messy reality of dialog environments. Qualitative examples illustrating these failure modes are provided in D.7. Future work will explore handcrafted and learned hybrid state representations, more robust uncertainty calibration for System 2, and adaptive or continuous exploration bonuses to mitigate these issues.
@@ 6 Conclusion
DyBBT introduces a principled, cognitively grounded framework for dialog policy learning that dynamically balances exploration and exploitation through a bandit inspired meta-controller operating over a structured cognitive state space. By formalizing dialog affordances, phasic progression, user uncertainty, and slot dependency, our approach enables adaptive, context aware switching between fast intuitive responses and deliberate reasoning. Extensive experiments demonstrate state-of-the-art performance across single and multi domain benchmarks, with human evaluations confirming superior decision quality and alignment with expert judgment. DyBBT offers a scalable, efficient, and interpretable alternative to static or population based methods, bridging cognitive theory with practical dialog optimization. Future work will focus on end-to-end learning of cognitive representations and extending the framework to more complex, interactive settings.

### v2

@@ 4.1 Experimental Setup
Datasets. We evaluate DyBBT on the Microsoft Dialog Challenge (Li et al., 2018) for single-domain tasks, and MultiWOZ 2.1 (Eric et al., 2020) for multi-domain tasks. Both are widely adopted in prior work. See Appendix B.1 for statistics.
Baselines. We compare DyBBT with a comprehensive set of recent and competitive baselines to ensure a rigorous evaluation, including previous SOTA EIERL. Full details are provided in Appendix B.2.
Evaluation Metrics. For single-domain tasks: success rate, average turns, and reward (following EIERL (Zhao et al., 2025): +2t+2t for success, t-t for failure, 1-1 for every turn). For multi-domain: Inform, Success, Book rates, and Avg. Turns (formulas in Appendix B.3).
Implementation Details. Following EIERL for fair comparison, dialogs are capped at 30 (single-domain) and 40 (multi-domain) turns. Training runs for 500 epochs (single) and 10K epochs (multi). DyBBT uses the same Qwen3 (0.6B-8B) for both S1 and S2. Full details in Appendix B.5.
4.2 Main Results Analysis
@@
Table 1: Evaluation results for all agents across the three single-domain datasets are provided, with the highest value in each metric column highlighted in bold. Epochs (50, 250, 500) represent early, mid, and post convergence training stages. Baselines sourced from Zhao et al. (2025).
@@
4.2.1 Performance on single-domain Tasks
The results in Table 1 show that DyBBT achieves strong performance across all three single-domain tasks. DyBBT’s cognitive state representation enables more efficient policy learning, leading to higher success rates with fewer turns compared to baselines. This advantage is especially pronounced in complex domains like Taxi, where slot dependencies create challenging exploration landscapes that DyBBT navigates effectively through its principled switching mechanism.
4.2.2 Performance on multi-domain Tasks
Appendix Table 6 presents results on the challenging MultiWOZ dataset. DyBBT maintains strong performance, whereas EIERL’s success rate drops significantly, revealing scalability limits of its population based approach. DyBBT-8B outperforms AutoTOD and ProTOD, and using GPT-4 as S2 achieves SOTA results, demonstrating that DyBBT matches strong LLM baselines with greater efficiency. This is enabled by the structured cognitive state and dual-system design, which provide domain-agnostic inductive bias without task-specific tuning. Cost-effectiveness analysis is discussed in Appendix E.7.
4.2.3 Training Efficiency and Convergence
Fig. 3 illustrates that DyBBT converges faster and achieves higher asymptotic performance than baselines across all domains, significantly outperforming EIERL as early as epoch 50. This accelerated learning stems from the meta-controller’s active guidance, which systematically targets under explored or uncertain regions in $\mathcal{C}$ instead of relying on random or high variance exploration. DyBBT also scales consistently with model size: success rates improve from 80.3% to 89.5% in the single-domain Movie task and from 78.2% to 84.1% in multi-domain settings when scaling from 0.6B to 8B. This indicates that the dual-system architecture effectively harnesses model ability. Coupled with efficient resource allocation via the meta-controller and Qwen3’s native switching mechanism (Appendix E.8), DyBBT demonstrates practical viability for real world deployment.
4.2.4 Key Advantages of DyBBT
The main results demonstrate that DyBBT achieves SOTA performance through the following key advantages: Dynamic Exploration Exploitation Balance: The bandit-inspired switching rule of the meta-controller enables DyBBT to allocate expensive S2 reasoning only when necessary, leading to highly efficient exploration. Scalability with Model Size: DyBBT benefits predictably from larger backbone models, making it well suited for future advances in LLM abilities. Strong Generalization: Consistent performance across both single- and multi-domain tasks shows that the cognitive state representation captures universal dialog dynamics. Computational Practicality: DyBBT maintains moderate computational overhead during both training and inference, unlike population based methods or full GPT-4.0 approaches.
@@ 4.3 Ablation Experiment
Table 2: Ablation study of DyBBT’s components on MultiWOZ. Results underscore the necessity of the meta-controller and the structured cognitive state representation for optimal performance.
[…]
@@
Ablation results are shown in Table 2, and detailed settings are in Appendix E.2. The results reveal that: Meta-Controller is crucial. Removing it causes the most severe performance degradation, confirming its essential role in dynamically orchestrating the exploration-exploitation trade-off. Both conditions are necessary but asymmetric: Removing Condition 1 (EC) eliminates the bandit inspired exploration bonus from Eq. 2, while removing Condition 2 (CC) disables the aleatoric uncertainty safeguard, a distinction rooted in Bayesian RL theory (Dearden et al., 1998). Removing the confidence condition (CC) causes a more substantial performance drop than removing the exploration condition (EC), validating our hybrid design. This indicates that mitigating S1’s overconfidence is slightly more critical than targeted exploration for robust performance. In depth error analysis (Appendix E.3) reveals that CC primarily prevents catastrophic failures in states with high cognitive uncertainty. Cognitive State design is vital. Replacing it with the raw belief state causes catastrophic performance collapse, confirming the necessity of our low dimensional, interpretable representation. While the learned alternative performs reasonably well, it still underperforms our hand-designed features, justifying our cognitively inspired approach. All state dimensions contribute meaningfully. Removing any single dimension causes noticeable performance degradation, with dialog progress ($d_{t}$) being the most impactful individual component, followed by user uncertainty ($u_{t}$) and slot dependency ($\rho_{t}$). Knowledge Distillation enables continuous improvement. Disabling it reduces final performance, confirming its role in facilitating long term efficiency gains through systematic learning from S2’s demonstrations.
4.4 Human and Real World Evaluation
We conducted controlled human evaluations and real world user experiments to validate DyBBT’s practical efficacy beyond automated metrics.
@@ 4.4 Human and Real World Evaluation
Human Evaluation. Following the protocol in Appendix C, 10 NLP researchers evaluated 200 dialog states from MultiWOZ, comparing DyBBT against random switching and S1-only baselines. Annotators rated action appropriateness (5-point Likert scale) and judged whether invoking S2 was justified. DyBBT’s actions were rated as more appropriate than both baselines, and its decisions to invoke S2 aligned significantly better with human judgment than random switching, confirming that our meta-controller effectively identifies when deliberation is warranted, an affordance often missed by heuristic approaches.
Real World User Experiments. As detailed in Appendix D, 30 volunteers completed multi-domain dialog tasks. DyBBT achieved the highest task success rate and user satisfaction, demonstrating that its cognitive state representation $\mathcal{C}$ generalizes effectively beyond simulated environments. Case studies further showed that DyBBT successfully handles challenging scenarios, such as mid-dialog intent shifts and vague user expressions through adaptive S2 invocation.
Summary. These results provide converging evidence that DyBBT’s meta-controller translates cognitive affordances into a dynamic exploration exploitation balance, enabling robust performance in both controlled and real world settings.
@@ 5 Analysis
Our experimental results demonstrate that DyBBT achieves state-of-the-art performance on multiple benchmarks. In this section, we analyze the underlying mechanisms that enable DyBBT’s effectiveness, providing insights into why and how our framework works.
5.1 Analysis of Cognitive State Space
@@
The cognitive state space $\mathcal{C}$ enables bandit style exploration in dialog POMDPs by providing a low dimensional and interpretable representation of dialog dynamics. To validate its structure, we analyze the visitation frequency across discretized regions of $\mathcal{C}$ during training (Fig. 5). The visitation heatmap reveals a structured, non-uniform pattern, confirming that exploration is guided by cognitive affordances: In early dialog phases ($d_{t}\in[0.0,0.2]$), the meta-controller broadly explores across high user uncertainty ($u_{t}$) for information gathering. In mid-phase ($d_{t}\in[0.4,0.6]$), visitation concentrates in medium-to-high $u_{t}$ regions to resolve ambiguities. In late phase ($d_{t}>0.8$), activity shifts to low $u_{t}$ states, focusing on exploitation to complete tasks.
This phase dependent targeting demonstrates that $\mathcal{C}$ effectively captures dialog progression and uncertainty, allowing the meta-controller to allocate exploration efficiently. The compactness and interpretability of $\mathcal{C}$ make principled exploration feasible in high dimensional dialog state spaces.
5.2 Adaptive Balancing and Continuous Improvement
The meta-controller’s hybrid triggering mechanism robustly addresses the exploration exploitation dilemma by responding to complementary forms of uncertainty. Condition 1 tackles epistemic uncertainty (lack of environmental knowledge), systematically exploring novel cognitive states, while Condition 2 addresses aleatoric uncertainty (inherent stochasticity or model limitations), acting as a consistent safety net against over reliance on a potentially flawed S1. Analysis of 10,000 dialog turns (Appendix Fig. 5) reveals their complementary temporal patterns: the exploration condition dominates early in training and for novel states, enabling systematic coverage, whereas the confidence condition provides ongoing robustness throughout training.
This dual trigger design naturally evolves with training progress. Initially, frequent S2 invocations provide guided exploration and high quality demonstrations. As S1 improves via knowledge distillation from these demonstrations, the meta-controller automatically reduces S2 usage, transitioning from guided exploration toward autonomous operation. This virtuous cycle enables continuous policy improvement without additional environment interactions. The distillation effectiveness is evidenced by monotonic improvement in S1 performance and corresponding reduction in S2 invocation rate. (Appendix Fig. 7). The adaptive balancing directly embodies the framework’s ability to perceive and respond to dynamic dialog affordances, ensuring appropriate cognitive resource allocation throughout the learning process.
@@ 5.3 Theoretical Intuitions and Empirical Alignment
Our theoretical analysis, though based on simplifying assumptions, is pragmatically validated by empirical results: Sublinear Regret as Validation of Core Assumptions. The empirical cumulative regret (Fig. 7) exhibits $\sqrt{T}$-like growth. This sublinear trend is not merely observational; it provides indirect empirical support for our key theoretical assumptions: The Lipschitz continuity of the reward in $\mathcal{C}$ (Assumption 3.1), and the approximate structure of MDP over $\mathcal{C}$ (Assumption A.1). The alignment between theory and experiment suggests $\mathcal{C}$ effectively captures the latent structure enabling efficient exploration. Low Dimensional $\mathcal{C}$ Enables Practical Implementation. The consistent high performance of DyBBT using only a three dimensional cognitive state demonstrates that the essential features governing exploration (dialog progress, user uncertainty, slot dependency) can be distilled into a compact representation. This reduction in dimensionality is theoretically motivated by the dependence of the regret bound’s $\sqrt{\dim(\mathcal{C})}$ (Appendix A.2.2).
5.4 Failure Mode Analysis
DyBBT exhibits three concrete failure modes empirically validated in Appendices E.9 and E.10. First, reliance on handcrafted cognitive state fidelity can lead to misrepresentation of complex dialog dynamics, causing the meta-controller to misjudge S2 invocation and resulting in underexploration or computational waste. Second, dependency on high quality S2 demonstrations introduces risk; errors in reasoning or self-evaluation can propagate to S1 via knowledge distillation, causing subtle policy corruption. Third, heuristic quantization of $\mathcal{C}$ into a fixed number of bins masks critical state variations, treating strategically distinct states identically and reducing exploration efficacy. Qualitative case studies illustrate how failures arise from unrepresented dialog nuances and how successful interventions align with human judgment, and reveals these failures affect only 5.2% of dialogs, primarily in edge cases with abrupt intent shifts or complex dependencies, while built-in safeguards provide substantial mitigation.
@@ 6 Conclusion
DyBBT presents a novel dialog policy learning framework that dynamically balances exploration and exploitation through a bandit inspired meta-controller grounded in a structured cognitive state space. By formalizing dialog affordances into interpretable dimensions: phasic progress, user uncertainty, and slot dependency, our method enables adaptive switching between fast intuitive responses and deliberate reasoning. Extensive experiments across single- and multi-domain benchmarks demonstrate SOTA performance in success rate, efficiency and generalization, with human evaluations confirming superior decision alignment. Future work will explore end-to-end learning of cognitive representations and extend the framework to more complex interactive settings.
Limitations
While DyBBT demonstrates strong empirical performance, its reliance on hand designed cognitive state representations, may not capture all nuances of highly complex or novel dialog dynamics. This points to a natural direction for future work: exploring end-to-end learning of cognitive representations that can adaptively refine the state space from interaction data, thereby extending the framework’s applicability to more diverse and intricate interactive settings.
Ethics Statement
This work presents a dialog policy learning framework evaluated on publicly available benchmark datasets (MS Dialog and MultiWOZ) and their use complies with the consent agreements established during their original release. These datasets have been previously anonymized and do not contain personally identifiable information or offensive content. Our research does not involve human subjects beyond the use of standard datasets, and all experiments are conducted through simulated user interactions. In the human evaluation and real user experiments, participants were volunteers proficient in English and knowledgeable in NLP, recruited from academic networks in Europe, the United States, and China. No monetary compensation was provided. Informed consent was obtained from all participants, and the study protocol followed established ethical guidelines for non-interventional research. The proposed methodology focuses on improving the efficiency of task-oriented dialog systems, with potential positive societal impacts through enhanced human computer interaction. We are unaware of any specific ethical concerns or negative social impacts directly arising from this work.
We utilized DeepSeek V3.2 for translation assistance and grammatical refinement of certain textual passages, and employed Qwen3-Code to aid in debugging and optimizing portions of the experimental code. These LLMs served solely as support tools for improving linguistic clarity and technical implementation. They played no role in the conceptualization of the research. The authors assume full responsibility for all aspects of the work, including the accuracy and integrity of all generated and modified content, and affirm that appropriate measures have been taken to prevent plagiarism and other forms of scientific misconduct.